Jacky is twice as old as lacy. In three years the sum is f their ages will be 54. Find their present age
Found 2 solutions by ewatrrr, addingup:
Answer by ewatrrr(24785)

x Lacy's age, Jacky 2x
Question States
(x+3) + (2x+3) = 54
3x+ 6 = 54
3x = 16
x = 16, Lacy's age. Jacky 32
and
in 3 years
19 + 35 = 54 CHECKS!

Answer by addingup(3677)

J is 2L, we can express this as:
J = 2L We'll use this value for J next
J+3+L+3 = 54 Substitute for J
2L+3+L+3 = 54
3L+6 = 54
3L = 48
L = 48/3 = 16 Lacy is 16
and
J = 2L = 2(16) = 32 Jacky is 32